# CycleGAN-Driven Transfer Learning for Electronics Response Emulation in High-Purity Germanium Detectors

### v1

@@ 4.4 Discussion
Our use case of the Cycle-GAN technique allows for quantitative studies of its performance using well-studied and physically meaningful parameters. In addition to these studies, we can also conduct a qualitative evaluation. Fig. 5 shows that the data-like pulse shapes resulting from applying CPU-Net include many of the more-subtle features seen in the data, such as overshoot at the end of each rising edge and the impact of multiple RC decay constants. This suggests that as novel PSD parameters and techniques are developed, simulated pulses with CPU-Net translation applied to them will serve as an accurate model for true data pulses.
CPU-Net learns electronics-response features that are present in every pulse, but it deliberately does not add any physics effects absent from the input simulation to the pulses. For instance, the narrow current-amplitude peak of DEP events in simulations suggests that our input simulations are generating single-site event pulses that are too uniform across the detector, possibly due to missing microphysics effects like charge cloud shape distortion and bulk charge trapping. The CPU-Net method does not obscure these disagreements, allowing it to be used in studies of how to improve the underlying HPGe detector simulations.
Although our main focus in this work was to translate simulations to resemble measured data, CPU-Net’s bidirectional nature also makes it a powerful tool for denoising real detector pulses. By transforming noisy detector signals into cleaner, simulation-like pulses, CPU-Net can help isolate and identify subtle features in the event topology. This capability could ultimately improve spatial reconstruction and improve our understanding of particle interactions within the detector.
@@ 6 Acknowledgment
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Workforce Development for Teachers and Scientists, Office of Science Graduate Student Research (SCGSR) program. The SCGSR program is administered by the Oak Ridge Institute for Science and Education for the DOE under contract number DE-SC0014664.

### v2

@@ 4.4 Discussion
Our use case of the Cycle-GAN technique allows for quantitative studies of its performance using well-studied and physically meaningful parameters. In addition to these studies, we can also conduct a qualitative evaluation. Figures 6 and 7 show that the data-like pulse shapes resulting from applying CPU-Net include many of the more-subtle features seen in the data, such as overshoot at the end of each rising edge and the impact of multiple RC decay constants. This suggests that as novel PSD parameters and techniques are developed, simulated pulses with CPU-Net translation applied to them will serve as an accurate model for true data pulses.
CPU-Net seeks to determine the transformation $\Lambda$ that will take each simulated pulse to the equivalent realistic detector pulse; its structure is, by design, blind to ensemble-level features like the width of peaks in the $T_{drift}$ and $I_{max}$ distributions, which emerge from variations in these parameters among pulses rather than an overall effect applied to all pulses. Effects like the overall too-narrow peaks observed in the simulations used for this work have been connected to unmodeled physics effects in the siggen software package. Tests employing a more-realistic charge cloud model, for example, show broader peaks in these distributions, improving agreement with the data. These ensemble-level discrepancies are preserved by CPU-Net, so the overall network performance on the IoU metrics has the potential to improve with the use of more realistic pulse simulations.
4.5 Uncertainty Evaluation
To measure the uncertainty associated with the network initialization process, we trained 500 trials of CPU-Net using different random seeds. The IoU over $I_{max}$ and $T_{drift}$ are then calculated for each trial over the same SEP validation dataset of 12,000 pulses. Figure 10 shows the distribution of IoU values for the current amplitude and drift time distributions. The optimized network results, also shown in this figure, are biased towards higher values of the IoU metrics because hyperparameter tuning was performed using a fixed seed value.
The multi-peaked structure in these distributions appears to be a consequence of stochastic training rather than a physical effect. Each point in those histograms is an independent training run with a different random seed, and the adversarial training appears to fall into a few distinct convergence modes. These modes can be interpreted as different local minima in the $\sim 14.2$M parameter space of the full model (two generators and two discriminators).
Since many of the hyperparameters are tuned to encourage network convergence when starting from the initial weights associated with that specific seed, varying the seed value can lead to failed GAN training, where one component of the adversarial network overpowers another. This indicates that beginning from pre-trained network weights may be a helpful strategy in scaling CPU-Net so it can be applied to multiple detectors and varying run conditions.
Figure 10: IoU distribution of 500 CPU-Net trials, varying the weight initialization, over maximal current amplitude $I_{max}$ (left) and drift time $T_{drift}$ (right). The distributions quantify the spread in performance arising from stochastic optimization and initialization, providing an estimate of the uncertainty of CPU-Net performance for fixed hyperparameters. The multi-peaked structure reflects that the adversarial training tends to converge into a small number of distinct modes in local minima in the parameter space, rather than a single unique solution. The low IoU tail corresponds to runs in which the GAN training becomes unstable, for instance when the generator or discriminator overpowers its counterpart, which lead to poorer alignment between translated and detector distributions. The dashed red lines indicate the IoU values of the optimized model used in the main analysis, which was obtained during hyperparameter tuning with a fixed random seed.
@@ 5.1 Summary of Results
5.2 Future Work
Although our main focus in this work was to translate simulations to resemble measured data, CPU-Net’s bidirectional nature also offers the potential for denoising real detector pulses. By transforming noisy detector signals into cleaner, simulation-like pulses, CPU-Net can help isolate and identify subtle features in the event topology. This capability could ultimately improve spatial reconstruction and improve our understanding of particle interactions within the detector.
HPGe detector measurements using the Compton scanning technique provide position-tagged, paired waveforms using a Compton-camera setup [Abt_2022odr]. Using data from such a system would enable direct quantitative per-pulse evaluation of CPU-Net. We leave this dedicated study for future work. Future studies will also study how CPU-Net generalizes to other detectors, geometries, and changing experimental conditions.
@@ 6 Acknowledgment
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Workforce Development for Teachers and Scientists, Office of Science Graduate Student Research (SCGSR) program. The SCGSR program is administered by the Oak Ridge Institute for Science and Education for the DOE under contract number DE‐SC0014664.